# Utilizing Contextual Clues and Role Correlations for Enhancing Document-level Event Argument Extraction

### v2

@@ 5.1.4 Hyperparameter Setting
EACE zhou2024eace - - 50.1 56.3
RKDE hu2023role - - 50.3 55.1
Span-based
Two-Step zhang2020two 38.9 46.4 40.1 47.7 -
[…]
${\text{TSAR}}_{{\text{base}}}$ xu2022two 45.23 51.70 48.06 55.04 -
${\text{TSAR}}_{{\text{large}}}$ xu2022two 49.23 56.76 51.18 58.53 -
${\text{SCPRG}}_{{\text{base}}}$ liu-etal-2023-enhancing-document 46.06 52.75 48.18 55.83 -
${\text{SCPRG}}_{{\text{large}}}$ liu-etal-2023-enhancing-document 50.23 57.22 51.88 59.06 -
Our Methods
[…]
${\text{CARLG}}_{{\text{base}}}$ 46.83 53.42 49.81 56.21 55.91
${\text{CARLG}}_{{\text{large}}}$ 50.64 57.71 52.82 59.79 56.71
Span-based+Ours
Two-Step+Ours 43.42 50.58 45.27 52.79 50.34
${\text{TSAR}}_{{\text{base}}}$+Ours 46.52 53.28 49.23 56.07 53.22
${\text{TSAR}}_{{\text{large}}}$+Ours 50.61 57.96 52.52 59.77 57.03
5.2 Main Results
5.2.1 Performance of our enhanced baseline
[…]
5.2.2 Performance of CARLG
[…]
Table 4: Main results of WikiEvents and MLEE datasets. * means we rerun their code based on their experimental settings (hyperparameter tuning is conducted when necessary). The highest scores are highlighted in bold, while the second-highest scores are underlined.
WikiEvents MLEE
Method Arg IF Arg CF Arg IF Arg CF
Head F1 Coref F1 Head F1 Coref F1 Span F1 Span F1
Tag-based
BERT-CRF ebner2020multi 69.83 72.24 54.48 56.72 - -
Generation-based
BERT-QA du-cardie-2020-event 61.05 64.59 56.16 59.36 - -
BERT-QA-Doc du-cardie-2020-event 39.15 51.25 34.77 45.96 - -
EEQA du-cardie-2020-event - - 56.9 - 68.4 66.7
[…]
${\text{EA}^{2}\text{E}}_{{\text{large}}}$ zengetal2022ea2e 74.62 75.77 68.61 69.70 - -
EACE zhou2024eace - 71.1 - 66.2 - -
RKDE hu2023role - 69.1 - 63.8 - -
Span-based
Two-Step zhang2020two 70.32* 72.27* 55.11* 57.12* 70.64* 68.74*
${\text{TSAR}}_{{\text{base}}}$ xu2022two 75.52 73.17 68.11 66.31 72.21* 71.03*
${\text{TSAR}}_{{\text{large}}}$ xu2022two 76.62 75.52 69.70 68.79 73.92* 72.71*
${\text{SCPRG}}_{{\text{large}}}$ liu-etal-2023-enhancing-document 77.52 76.01 70.11 69.25 73.88* 73.21*
Our Methods
[…]
${\text{CARLG}}_{{\text{base}}}$ 76.22 74.97 69.03 68.53 73.26 71.23
${\text{CARLG}}_{{\text{large}}}$ 77.22 76.03 70.92 70.11 75.90 73.71
Span-based+Ours
Two-Step+Ours 73.20 72.96 60.90 61.04 72.42 70.11
${\text{TSAR}}_{{\text{base}}}$+Ours 76.43 75.10 68.81 68.22 73.13 71.12
${\text{TSAR}}_{{\text{large}}}$+Ours 77.46 76.10 70.88 70.28 74.98 73.48
5.2.3 Performance of Span-based Methods equipped with CARLG
In order to evaluate the compatibility of our CCA and RLIG modules, we equip these two modules on some classical span-based methods and conduct experiments on three datasets. As is shown in Table 3 and Table 4, both Two-Step and TSAR exhibit improved performance when equipped with CCA and RLIG modules. Two-Step yields an improvement of +5.17/+5.09 Span F1/Head F1 for Arg CF on RAMS, +2.88/+5.79 Head F1 for Arg IF/Arg CF on WikiEvents, and +1.78/+1.37 Span F1 for Arg IF/Arg CF on MLEE. ${\text{TSAR}}_{{\text{large}}}$ achieves notable improvements in Span F1/Head F1 for Arg CF on RAMS (+1.34/+1.24), Head F1 for Arg IF/Arg CF on WikiEvents (+0.84/+1.18), and Span F1 for Arg IF/Arg CF on MLEE (+1.06/+0.77). It is worth noting that the improvement of TSAR after incorporating CCA and RLIG modules is relatively modest. This could be attributed to the fact that the AMR graph already contains contextual semantic information, which may mitigate the need for additional contextual clue integration. These results illustrate that our two modules can be easily applied to other span-based methods with transformer encoders and provide a significant performance boost in document-level EAE tasks, which are compact and transplantable.
[…]
Method Dev Test
Span F1 Head F1 Span F1 Head F1
[…]
 -Boundary Loss 42.18 50.28 42.33 51.68
 -Focal Loss 44.97 51.36 45.83 53.03
 -FIS 43.27 50.64 45.26 52.11
Table 6: Ablation Study on RAMS and WikiEvents for CARLG. Span, Head and Coref refer to Span F1, Head F1 and Coref F1 scores respectively.
RAMS WikiEvents
Method Params Dev Test Arg IF Arg CF
Span Head Span Head Head Coref Head Coref
${\text{CARLG}}_{{\text{base}}}$ 120.87M 46.83 53.42 49.81 56.21 76.22 74.97 69.03 68.53
-CCA 120.28M 45.78 52.67 47.43 54.57 74.57 73.86 67.44 66.87
-RLIG 118.78M 45.27 52.36 47.91 54.52 75.59 73.83 68.26 67.49
-both 118.19M 45.07 51.59 45.76 53.16 73.92 73.22 66.90 65.98
${\text{CARLG}}_{{\text{large}}}$ 374.70M 50.64 57.71 52.82 59.71 77.22 76.03 70.92 70.11
-CCA 373.65M 49.68 56.46 50.84 57.9 75.68 73.82 69.79 68.65
 -RLIG 371.68M 49.96 57.32 51.28 58.46 76.30 74.08 69.87 68.96
-both 370.63M 48.33 54.04 47.52 55.61 75.45 73.55 68.63 67.30
5.3 Ablation Study
In order to effectively showcase the effectiveness of our components, we carry out an ablation study on CARLG using the RAMS and WikiEvents datasets, as depicted in Table 6. First, when we remove the Contextual Clues Aggregation (CCA) module, we observe a decrease in Span F1 and Head F1 scores of ${\text{CARLG}}_{{\text{base}}}$/ ${\text{CARLG}}_{{\text{large}}}$ by 2.83/1.98 and 1.64/1.81 on the test set, respectively, which indicates that our CCA module plays a crucial role in capturing essential contextual clues for document-level EAE.
Moreover, when removing the Role-based Latent Information Guidance (RLIG) module [^3] We additionally remove the respective role tokens that were included in the input sequence., we observe a significant decrease in performance for ${\text{CARLG}}_{{\text{base}}}$/ ${\text{CARLG}}_{{\text{large}}}$, with a drop of 1.9/1.54 in Span F1 and 1.69/1.07 in Head F1 on the RAMS test set. This demonstrates the effect of the RLIG module in capturing meaningful semantic correlations among roles, improving the performance of document-level EAE. Moreover, when both the CCA and RLIG modules are removed, the decline in performance is even more noticeable in comparison to excluding a single module. This illustrates the synergistic enhancement achieved by the collaboration of these two modules.
[…]
Figure 7: Context weights for a sample from RAMS are displayed, focusing on the span-trigger pair (Islamic State, wounded). Various color intensities are used to depict the attention weights.
@@ 5.4 Analysis of Contextual Attentive Weights
To evaluate the validity of our Contextual Clues Aggregation (CCA) module in capturing beneficial contextual clues for candidate argument spans, we present a visualization of the contextual weights $\textbf{p}^{c}_{i:j}$, as detailed in Eq.12, using an instance from the RAMS dataset. As shown in Figure 7, the CCA module assigns high weights to non-argument contextual clues such as attack, responsibility, and terrorist attack, which are highly relevant to the span-trigger pair (Islamic State, wounded). Notably, our Contextual Clue Aggregation (CCA) module also grants relatively high attentive weights to some words in other arguments, including explosive, Dozens, and Kabul. This implies that these argument words offer crucial information for predicting the role of Islamic State. This visualization illustrates how our Contextual Clue Aggregation (CCA) module effectively identifies relevant contextual clues for candidate spans and models the interaction of information among interconnected arguments within an event.
@@
Additionally, our investigation extends to examining the attention weights associated with various span-trigger pairings within a single event. In Figure 8, we display a heat map created from the attention weights of 30 randomly chosen candidate spans in an event, with respect to their contextual relevance. The heat map reveals that different candidate arguments pay attention to different contextual information, suggesting that our Contextual Clues Aggregation (CCA) module is capable of dynamically identifying and utilizing pertinent contextual clues specific to each candidate argument span. This adaptability highlights the module’s efficiency in tailoring its focus to suit the unique contextual requirements of each argument span within an event.
Figure 9: The visualization of cosine similarity between role representations of two examples from RAMS in sub-figure (a) and MLEE dataset in sub-figure(b) respectively.
Figure 10: Here’s an instance of a t-SNE visualization sourced from the RAMS dataset. The representations of roles and arguments originate from five distinct documents. In sub-figure (a), we utilize average pooling representations encoded by BERT, while in (b), we visualize the representations fused with latent role embeddings.
@@ 5.5 Analysis of Role Correlations
To verify the capability of our model in capturing semantic correlations between roles, we visualize the cosine similarity among latent role representations from events within the RAMS and MLEE datasets. Figure 9 presents the results, where roles such as origin and destination, attacker and target, ToLoc and FromLoc exhibit similar representations, indicating their semantic similarity. This visualization validates our model’s efficacy in capturing semantic correlations among roles.
Additionally, to substantiate the effective guidance provided by role representations, we utilize t-SNE visualization JMLR:v9:vandermaaten08a to display arguments corresponding to two distinct roles that appear together in five diverse documents, alongside their respective latent role embeddings. Figure 10a illustrates that arguments associated with the same role but in distinct documents are dispersed throughout the embedding space. This dispersion is attributed to variations in the target events and their contextual settings. However, after integrating latent role embeddings, the representations of arguments linked to identical roles, such as victim or place, cluster more closely. This shows that the Role-based Latent Information Guidance (RLIG) module effectively provides crucial role-specific information. The visualization demonstrates that the RLIG module enhances the alignment of arguments within the same role across different contexts, thereby affirming its utility in providing insightful guidance of roles.
@@ 5.6 Complexity and Efficiency
Our CARLG method combines computational efficiency with low complexity while both CCA and RLIG introduce few parameters. As shown in Table 6, for ${\text{CARLG}}_{{\text{large}}}$, CCA utilizes the pre-trained encoder’s attention heads and performs multiplication and normalization operations, adding only about 0.28% new parameters. RLIG introduces approximately 0.81% new parameters totally. This makes our parameter quantity approximately equivalent to the transformer-based encoder plus a MLP classification layer.
Table 7 reports the efficiency comparison of various span-based models on MLEE dataset. Our ${\text{CARLG}}_{{\text{base}}}$ is approximately 2 times faster than ${\text{TSAR}}_{{\text{base}}}$ and can save up to 30.9% of inference time with smaller model size. Additionally, compared to Two-Step, our CARLG only increases 15.9% inference time and 8% model parameters. Such high efficiency and performance makes CARLG a promising choice for practical applications of document-level EAE.
Table 7: Time cost (in seconds) of different span-based models on MLEE dataset. Experiments are run for one epoch on one Tesla A100 GPU.
@@
Two-Step 111.96M 229.68 55.29
${\text{TSAR}}_{{\text{base}}}$ 142.03M 465.84 92.79
${\text{CARLG}}_{{\text{base}}}$ 120.87M 280.08 64.08
@@ 5.7 Error Analysis
Table 8: Error analysis on the RAMS test set. The best results are in bold font.
@@
${\text{TSAR}}_{{\text{base}}}$ 81 48 57 28 19 233
${\text{CARLG}}_{{\text{base}}}$ 79 45 55 29 15 223
@@
To conduct a more comprehensive examination of the errors made by various models and analyze the underlying reasons for these errors , we manually compared the predictions of 200 randomly selected examples from the RAMS test set against their golden annotations. During our analysis, we categorized the errors into five main categories following xu2022two . Wrong Span refers to the case where a specific role is assigned to a span that does not overlap with the corresponding golden span. We observe that this type of errors often occur when negative words such as ”not” are involved or when there are coreference spans related to the golden span. The category labeled Over-extract pertains to instances where the model predicts an argument role that does not genuinely appear in the document. Some of the extracted spans are either sub-strings of the golden spans (Partial) or exhibit overlaps with them (Overlap). The ”Partial” error category is also commonly observed when there is punctuation, such as a comma, within the golden span. Wrong Role refers to the model correctly identifies the golden span, but assigns the wrong argument role to it. We compare the errors of some span-based methods, that is, Two-Step and ${\text{TSAR}}_{{\text{base}}}$ and ${\text{CARLG}}_{{\text{base}}}$. As shown in Table 8, compared with Two-Step and ${\text{TSAR}}_{{\text{base}}}$, our CARLG model decreases the number of errors from 275 to 223 and effectively reduces Wrong Span errors, which can be attributed to the integration of contextual clues and latent role guidance.
@@
CARLG (Ours) 56.71 52.82
@@ 5.8 Comparison with Large Language Models
Large language models (LLMs) have become a focal point of interest among researchers, given their remarkable versatility across a diverse range of Information Extraction (IE) tasks xu2023large ; li2023semi ; zhang2024ultra . In this paper, we undertake a comprehensive comparison with some large language models on the document-level EAE task. Due to the high cost associated with using GPT-4, its evaluation is limited solely to the RAMS dataset. The experimental results are detailed in Table 9. We utilize the prompts defined in the work zhou2023heuristics and experiments are conducted utilizing three prominent large language models: text-davinci-003 ouyang2022training , gpt-3.5-turbo ouyang2022training and GPT-4 achiam2023gpt , all of which are accessed via OpenAI’s public APIs  [^4] https://openai.com/api/. As illustrated in Table 9, LLMs exhibit a notable performance disparity compared to supervised learning models in the EAE task. Furthermore, the operational costs associated with large models are inherently substantial. In contrast, our approach demonstrates superior efficiency, cost-effectiveness, and overall effectiveness in the document-level EAE task when compared to LLMs.
@@ 5.9 Case Study
In order to explore how the CCA and RLIG modules work in actual cases, we conduct the case study for our CARLG. As shown in Figure 11, equipped with CCA and RLIG modules, our CARLG can correctly predict all arguments, while TSAR fails to predict arguments attack. When removing the CCA module, our CARLG fails to predict the argument Talha Hesham Abdo, indicting the CCA module contributes to the argument extraction by providing essential contextual clues. Additionally, when the RLIG module is removed, it fails to predict the argument for the crim role and mispredicts the argument for the jailer role. This can be attributed to that the RLIG module effectively captures the semantic correlations among the roles jailer, detainee, and crim and provides valuable information guidance, allowing the model to make more accurate predictions.

### v4

@@ 6.3 Ablation Study
Figure 7: Attentive weights for an event instance from RAMS dataset. The event is life.die.death-caused-by-violent-events, focusing on the argument government and trigger bombarding). Various color intensities are used to depict the attention weights. We conduct visualization experiments on $\text{CARLG}_{\text{span}}$ and $\text{CARLG}_{\text{prompt}}$, selecting TSAR and PAIE as the backbone methods, respectively.
@@ 6.4 Analysis of Contextual Attentive Weights
To evaluate the validity of our Contextual Clues Aggregation (CCA) module in capturing beneficial contextual clues for candidate argument spans, we present a visualization of the contextual weights $\textbf{p}^{C}_{a_{i}}$, as detailed in Eq.1, using an instance from the RAMS dataset. We conduct visualization experiments on $\text{CARLG}_{\text{span}}$ and $\text{CARLG}_{\text{prompt}}$, selecting TSAR and PAIE as the backbone methods, respectively. As shown in Figure 7, the CCA module assigns high weights to non-argument contextual clues such as surrender, starving, and shooting, which are highly relevant to the argument-trigger pair (government, bombarding). Notably, our Contextual Clue Aggregation (CCA) module also grants relatively high attentive weights to some words in other arguments, including shelling, people. This implies that these argument words offer crucial information for predicting the role of government. Additionally, we find that for $\text{CARLG}_{\text{prompt}}$, the CCA module also learns high attention towards prompt words such as killed and die, indicating that the CCA module assists the model in better understanding the prompt. This visualization illustrates how our CCA module effectively identifies relevant contextual clues for target arguments and models the interaction of information among interconnected arguments within an event.
@@
Additionally, our investigation extends to examining the attention weights associated with various argument-trigger pairings within a single event. In Figure 8, we display a heat map created from the attention weights of 30 randomly chosen candidate spans in an event, with respect to their contextual relevance [^5] We conduct this experiment on span-based EAE methods TSAR (Xu et al., 2022b) equipped with our $\text{CARLG}_{\text{span}}$. The heat map reveals that different candidate arguments pay attention to different contextual information, suggesting that our Contextual Clues Aggregation (CCA) module is capable of dynamically identifying and utilizing pertinent contextual clues specific to each candidate argument span. This adaptability highlights the module’s efficiency in tailoring its focus to suit the unique contextual requirements of each argument span within an event.
(a) The similarity for $\text{CARLG}_{\text{span}}$ in RAMS
(b) The similarity for $\text{CARLG}_{\text{prompt}}$ in RAMS
(c) The similarity for $\text{CARLG}_{\text{span}}$ in MLEE
(d) The similarity for $\text{CARLG}_{\text{prompt}}$ in MLEE
Figure 9: The visualization of cosine similarity between role representations of two examples from RAMS and MLEE datasets, respectively.
Figure 10: Here’s an instance of a t-SNE visualization sourced from the RAMS dataset. We conduct this experiment using PAIE equipped with our $\text{CARLG}_{\text{prompt}}$ framawork. The representations of roles and arguments originate from five distinct documents. In sub-figure (a), we utilize average pooling representations encoded by BART (Lewis et al., 2019), while in (b), we visualize the representations fused with latent role embeddings.
@@ 6.5 Analysis of Role Correlations
To verify the capability of our model in capturing semantic correlations between roles, we visualize the cosine similarity among latent role representations. We select two events within the RAMS and MLEE datasets for $\text{CARLG}_{\text{span}}$ and $\text{CARLG}_{\text{prompt}}$ [^6] We select TSAR and PAIE as the backbone methods, respectively., respectively. Figure 9 presents the results, where roles such as origin and destination, attacker and target, ToLoc and FromLoc exhibit similar representations, indicating their semantic similarity. This visualization validates our model’s efficacy in capturing semantic correlations among roles.
Additionally, to substantiate the effective guidance provided by role representations, we utilize t-SNE visualization (van der Maaten & Hinton, 2008) to display arguments corresponding to two distinct roles that appear together in five diverse documents, alongside their respective latent role embeddings. Figure 10a illustrates that arguments associated with the same role but in distinct documents are dispersed throughout the embedding space [^7] We conduct this experiment using PAIE (Ma et al., 2022b) equipped with our $\text{CARLG}_{\text{prompt}}$ framawork. This dispersion is attributed to variations in the target events and their contextual settings. However, after integrating latent role embeddings, the representations of arguments linked to identical roles, such as victim or place, cluster more closely. This shows that the Role-based Latent Information Guidance (RLIG) module effectively provides crucial role-specific information. The visualization demonstrates that the RLIG module enhances the alignment of arguments within the same role across different contexts, thereby affirming its utility in providing insightful guidance of roles.
@@ 6.6 Complexity and Efficiency
Our CARLG method combines computational efficiency with low complexity while both CCA and RLIG introduce few parameters. As shown in Table 6, for CCA utilizes the attention heads of the pre-trained encoder and performs multiplication and normalization operations, adding less than 0.5% new parameters. RLIG introduces approximately 0.8% new parameters totally. This makes our parameter quantity approximately equivalent to the transformer-based encoder plus a MLP classification layer.
Figure 11: The parameter size, training time, and inference time of TSAR and PAIE before and after incorporating CARLG.
Table 7 reports the efficiency comparison of various span-based models on MLEE dataset. After integrating our $\text{CARLG}_{\text{span}}$ framework, ${\text{TSAR}}_{\text{C}}$ shows an increase of approximately 14.2% in training time and 11.4% in inference time. With the addition of $\text{CARLG}_{\text{prompt}}$, ${\text{PAIE}}_{\text{C}}$ shows an increase of about 7.4% in both training and inference time. The increased training and inference time associated with $\text{CARLG}_{\text{span}}$ is attributed to the exhaustive nature of the span-based EAE approaches, necessitating the enumeration of all possible spans. Consequently, the operations of the CCA and RLIG modules escalate in frequency as the quantity of candidate spans expands.
Table 7: The parameter size and time cost (in seconds) before and after integrating the PAIE and TSAR models with the CARLG framework on MLEE dataset. Experiments are run for one epoch on one Tesla A100 (40G) GPU.
@@
TSAR 392.44M 465.84 72.79
PAIE 405.12M 148.74 74.31
$\text{TSAR}_{\text{C}}$ 395.23M ($\uparrow$0.71%) 532.08 ($\uparrow$14.22%) 81.08 ($\uparrow$11.39%)
$\text{PAIE}_{\text{C}}$ 408.22M ($\uparrow$0.77%) 159.80 ($\uparrow$7.44%) 79.81 ($\uparrow$7.40%)
@@ 6.7 Error Analysis
Table 8: Error analysis on the RAMS test set. ${\text{TSAR}}_{\text{C}}$ and ${\text{PAIE}}_{\text{C}}$ represent the original TSAR and PAIE methods equipped with our CARLG framework. The best results are in bold font.
@@
TSAR 81 48 57 28 19 233
${\text{TSAR}}_{\text{C}}$ (Ours) 79 45 55 29 15 223
PAIE 61 46 59 34 25 225
${\text{PAIE}}_{{\text{C}}}$ (Ours) 59 45 59 29 18 210
@@
To conduct a more comprehensive examination of the errors made by various models and analyze the underlying reasons for these errors , we manually compared the predictions of 200 randomly selected examples from the RAMS test set against their golden annotations. During our analysis, we categorized the errors into five main categories following (Xu et al., 2022b). Wrong Span refers to the case where a specific role is assigned to a span that does not overlap with the corresponding golden span. We observe that this type of errors often occur when negative words such as “not” are involved or when there are coreference spans related to the golden span. The category labeled Over-extract pertains to instances where the model predicts an argument role that does not genuinely appear in the document. Some of the extracted spans are either sub-strings of the golden spans (Partial) or exhibit overlaps with them (Overlap). The “Partial" error category is also commonly observed when there is punctuation, such as a comma, within the golden span. Wrong Role refers to the model correctly identifies the golden span, but assigns the wrong argument role to it.
We compare the number of errors on TSAR and PAIE before and after adding the CARLG framework. As shown in Table 8, compared with Two-Step and TSAR, our ${\text{TSAR}}_{\text{C}}$ decreases the number of errors from 275 to 223 and effectively reduces Wrong Span errors, which can be attributed to the integration of contextual clues and latent role guidance. Meanwile, compared with the original PAIE method, the number of errors decreases from 225 to 210 after adding our $\text{CARLG}_{\text{prompt}}$, especially Wrong Role and Overlap errors. This fully demonstrates that our proposed RLIG module can capture semantic correlations between roles.
@@
$\text{TSAR}_{\text{C}}$ (Ours) 57.1 52.6
$\text{PAIE}_{\text{C}}$ (Ours) 59.1 54.4
@@ 6.8 Comparison with Large Language Models
Large language models (LLMs) have become a focal point of interest among researchers, given their remarkable versatility across a diverse range of Information Extraction (IE) tasks (Xu et al., 2023a; Li et al., 2023; Zhang et al., 2024). In this paper, we undertake a comprehensive comparison with some large language models on the document-level EAE task. Due to the high cost associated with using GPT-4, its evaluation is limited solely to the RAMS dataset. The experimental results are detailed in Table 9. We utilize the prompts defined in the work (Zhou et al., 2023) and experiments are conducted utilizing three prominent large language models: text-davinci-003 (Ouyang et al., 2022), gpt-3.5-turbo (Ouyang et al., 2022) and GPT-4 (Achiam et al., 2023), all of which are accessed via OpenAI’s public APIs  [^8] https://openai.com/api/. As illustrated in Table 9, LLMs exhibit a notable performance disparity compared to supervised learning models in the EAE task. Furthermore, the operational costs associated with large models are inherently substantial. In contrast, our approach demonstrates superior efficiency, cost-effectiveness, and overall effectiveness in the document-level EAE task when compared to LLMs.
@@ 6.9 Case Study
In order to explore how the CCA and RLIG modules work in actual cases, we conduct the case study for our CARLG [^9] We present a case study on $\text{CARLG}_{\text{span}}$, choosing TASR (Xu et al., 2022b) as the backbone model.. As shown in Figure 12, equipped with CCA and RLIG modules, our CARLG can correctly predict all arguments, while TSAR fails to predict arguments attack. When removing the CCA module, our CARLG fails to predict the argument Talha Hesham Abdo, indicting the CCA module contributes to the argument extraction by providing essential contextual clues. Additionally, when the RLIG module is removed, it fails to predict the argument for the crim role and mispredicts the argument for the jailer role. This can be attributed to that the RLIG module effectively captures the semantic correlations among the roles jailer, detainee, and crim and provides valuable information guidance, allowing the model to make more accurate predictions.